# Train-before-Test Harmonizes Language Model Rankings

### v1

@@ 3.2 Downstream Ranking Agreement
As depicted in Figure 2, direct evaluation shows only modest ranking agreement between the 24 benchmarks, with an average Kendall’s $\tau$ ranking correlation of 0.51. This lack of agreement across benchmarks complicates model assessment and makes it challenging to aggregate results into a meaningful overall ranking. In contrast, the train-before-test methodology leads to a substantial improvement in ranking agreement. Under this approach, 95% of model pairs show higher Kendall’s $\tau$ scores, with the average $\tau$ rising from 0.51 to 0.74. This improvement in ranking consistency is highly notable, as it indicates that models that excel on any given benchmark are also likely to excel in many others, thereby simplifying the model selection process for developers. Notably, benchmarks that appeared to be outliers under direct evaluation, such as MRPC and WiC, demonstrate much greater ranking consistency under train-before-test. For example, the average Kendall’s $tau$ between MRPC and all other benchmarks improves dramatically, from 0.23 to 0.73.
We further split all benchmarks into six categories (e.g., language understanding, math), see Table 1. For each category pair, we report in Figure 3 the intra-category average ranking correlations and inter-category average ranking correlations across all relevant benchmark pairs. Consistent with our previous findings, we observe reasonably poor ranking agreements across benchmark categories. While one might expect high intra-category agreement–after all, tasks within the same category tend to be relatively similar–direct evaluation results in low intra-category agreement in many cases. For example, the intra-category mean Kendall’s $\tau$ is 0.52 for language understanding and 0.56 for math. This further underscores the difficulty of selecting models based on direct evaluation. Even if the goal is to choose a model that excels not across all tasks but within a specific domain, the low intra-category agreement makes this decision challenging.
In contrast, train-before-test markedly boosts both intra- and inter-category consistency. For example, the intra-category mean Kendall’s $\tau$ for language understanding raises from 0.52 to 0.71, as well as from 0.56 to 0.82 for the math category. Moreover, agreement between categories is often nearly as high as agreement within categories. This suggests that models performing well in one domain tend to also perform well across other domains.
@@ 3.3 Perplexity Agreement
We now study the agreement between benchmark rankings and rankings based on perplexity on general domain corpora. To do so, we collect three corpora from Wikipedia, StackExchange, and arXiv, retaining only contents from 2025. More specifically, we collect 3,366 documents for Wiki, 6,001 for StackExchange and 44,384 documents for arXiv. These datasets are split into training, validation, and testing sets, in an 8:1:1 ratio.
@@
Figure 4: Ranking agreement between perplexity rankings and downstream benchmark rankings under direct evaluation (top) and train-before-test (bottom). Perplexity rankings are consistent with each other under both evaluation schemes, with an average Kendall’s $\tau$ of 0.76 and 0.78, respectively. However, for direct evaluation, agreement between perplexity rankings and downstream rankings is low, with an average Kendall’s $\tau$ of just 0.47. Fortunately, train-before-test results in much higher agreement between perplexity and downstream evaluations, increasing average Kendall’s $\tau$ to 0.73.
@@
Figure 5: Ranking agreement between perplexity rankings before fine-tuning and downstream benchmark rankings after fine-tuning, for base models (top) and instruction-tuned models (bottom). For base models, perplexity rankings before fine-tuning are consistent with downstream performance rankings after fine-tuning, with an average of Kendall’s $\tau$ 0.77. The same does not hold for instruction-tuned models, where the average Kendall’s $\tau$ between perplexity rankings and fine-tuning rankings is 0.5.
@@
We measure perplexity in bits per byte, using the lm-eval-harness library. We then construct models rankings based on the perplexity evaluations, and contrast the rankings with those of the downstream benchmarks considered in earlier sections. We exclude the four Gemma models from these results, as lm-eval-harness provides unreliable perplexity measurements for Gemma models due to its rolling window implementation. See Appendix B.2 for more details.
@@
The results are presented in Figure 4. In contrast to downstream tasks, perplexity rankings demonstrate strong agreement both for direct evaluation and train-before-test. Specifically, the average Kendall’s $\tau$ between the perplexity rankings is 0.76 for direct evaluation and 0.78 for train-before-test. We hypothesize that this reasonably strong agreement arises due to the smooth relationship between perplexity evaluations (brandfonbrener2024loss, ; mayilvahananllms, ).
@@
When comparing ranking agreement between perplexity evaluations and downstream benchmarks, we find that agreement is low under direct evaluation, with a mean Kendall’s $\tau$ of 0.47. This lack of agreement is concerning, as it signals a disconnect between the language modelling pre-training objective and downstream benchmark performance. Fortunately, we find that our train-before-test methodology improves ranking agreement substantially, with the mean Kendall’s $\tau$ ranking correlation between perplexity rankings and benchmark rankings rising to 0.73. This finding is reassuring: a light amount of fine-tuning on task data is sufficient to align the language modeling training objective with downstream performance. Moreover, we find that ranking agreement between perplexity and downstream evaluations is roughly similar to agreement across downstream evaluations. This suggests that, despite perplexity typically not being used for benchmarking purposes, it can be as effective a ranking metric as benchmark evaluations.
@@
Drawing inspiration from prior work (liu2023same, ; xia2023training, ; Gadre2024LanguageMS, ; benchmarkloss, ), we further examine the correlation between model rankings according to average perplexity across the three text corpora and average downstream performance across the 24 benchmarks. Gadre et al. (2024, (Gadre2024LanguageMS, )) show that, when models are trained on the same pretraining data, perplexity is well-correlated with aggregate benchmark performance. Our setup differs in that we consider a diverse set of model families, each trained on different pretraining data. Under direct evaluation, we find that the ranking correlation is modest, with a Kendall’s $\tau$ of only 0.56. We hypothesize that this relatively weak agreement is due to differences in pretraining data and instruction tuning, resulting in varying levels of exposure to benchmark tasks during training (dominguez2024training, ). Fortunately, when applying our train-before-test methodology, the ranking correlation between average perplexity and average downstream performance improves substantially, with Kendall’s $\tau$ increasing from 0.56 to 0.82.
@@
We additionally examine the agreement between perplexity prior to fine-tuning and downstream task performance after fine-tuning. That is, between direct evaluation perplexity rankings and train-before-test downstream performance rankings. We plot such ranking agreement in Figure 5, dividing models into base models and instruction-tuned models. We find that, for base models, perplexity prior to fine-tuning is a strong indicator of task performance after fine-tuning. This indicates that, for base models, direct evaluation of perplexity is already a reasonably reliable metric for ranking models. However, the same does not hold for instruction-tuned models. We find that instruction-tuning renders perplexity rankings unreliable, as ranking agreement is low across the board. This is to be expected: instruction fine-tuning tends to increase both benchmark performance and perplexity on general text corpora, thus clouding the relationship between perplexity and downstream evaluations. Fortunately, as shown earlier, train-before-test restores high ranking agreement between perplexity evaluations and downstream task performance.
@@
Figure 6: Explained variance ratios of the top five principal components of the benchmark score matrix, under direct evaluation (left) and train-before-test (right). Train-before-test substantially increases the amount of variance explained by the first principal component, from 69% to 85%
@@ 3.4 Low-Ranked Model Score Matrix
So far, we have shown that the train-before-test methodology significantly improves cross-benchmark ranking agreement. We now examine the implications of this finding by analyzing the resulting matrix of model scores, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$. We use Principal Component Analysis (PCA) to examine the structure of the matrix of model scores.
Figure 6 shows the explained variance ratios of the first five principal components. These results support previous findings that the score matrix is of low rank (Ruan2024ObservationalSL, ). Under direct evaluation, the first five components account for 91% of the total variance. A similar trend is observed for train-before-test scores, where the first five components explain 97% of the variance. Notably, under train-before-test, the first principal component (PC1) captures a significantly larger share of the variance: 85%, compared to 69% for direct evaluation.
Prior works interpret PC1 scores as an indication of general capability, with later principal components denoting domain-specific capabilities not captured by PC1 (e.g., reasoning ability, coding ability) (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ). Contrary to the popular notion that the performance of LLMs is governed by multiple axes of capability (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ), our results highlight the dominance of a single principal axis of general capability.
@@
It is of no surprise that PC1 positively correlates with pre-training compute, as shown in Figure 7222We only plot models whose number of training tokens is publicly available. See Table 4 for details., which have been identified as crucial to model performances (kaplan2020scaling, ). However, the exact interpretation of PC1 remains open to debate. It could be model scale (kaplan2020scaling, ), general capability (Ruan2024ObservationalSL, ), or another factor entirely. Nonetheless, it is evident that PC1 is strongly linked with model performance–particularly under the train-before-test methodology– and thus might serve as an effective target for guiding further model development. See detailed PC1 scores in Figure 11.
@@
We repeat our PCA analysis on the score matrix containing only Qwen models, depicted in Figure 8. Remarkably, we find that PC1 for train-before-test explains 93% of the variance, roughly as much as the variance explained by the top four principal components under direct evaluation. That is, whereas for direct evaluation the score matrix is low-rank; train-before-test renders the score matrix essentially rank one.
@@ 4 Discussion, limitations, and conclusion
Train-before-test harmonizes model rankings across numerous benchmarks we evaluated. Put differently, it endows model rankings with external validity they lack under direct evaluation. In turn, this greatly increases the utility of benchmarks for reliable model comparisons and model selection. To play devil’s advocate, however, one might ask why do we need ranking consistency if we can just look at the benchmark that best fits a downstream application? First, we show that even benchmarks that aim to capture the same task give contradictory rankings. Second, there will never be a perfect match between a benchmark and the downstream application. We must therefore always rely on some form of external validity, specifically, that the ranking transfers gracefully to other settings. Third, practitioners typically specialize any chosen model to their application. This makes direct evaluation numbers irrelevant, anyway.
Limitations.
Train-before-test requires that we fine-tune models on agreed upon task-specific data prior to evaluation. This certainly increases the cost of evaluation and might be too costly in some cases. On the other hand, it also points to interesting avenues for future work on how to implement train-before-test most efficiently. A second problem is that, unfortunately, many benchmarks no longer come with training data, making it more difficult to apply train-before-test. In light of our findings, we recommend that future benchmarks provide fine-tuning data for the benchmark. A third limitation is that some commercial model providers do not easily allow fine-tuning of their models. We contend that in this case the problem is with the model provider. There is clearly scientific value in creating an ecosystem of models that can be fine-tuned. Train-before-test evaluation creates additional incentives for making models easy to fine-tune.
Conclusion.
On the whole, we believe that train-before-test is a sound recommendation for the evaluation ecosystem that can help address some of its serious problems.
@@ A.1 Benchmark Selection
Language model evaluations can be broadly classified into likelihood-based and generation-based. Likelihood-based evaluations test for the likelihood of different completions given some input string; for example, different answer choices given a multiple-choice input question. Since the number of completions is usually small, likelihood-based evaluations are generally compute-efficient. Generation-based evaluations, in contrast, generate some output response given an input query. If responses tend to be long, then generation-based evaluations naturally become compute-intensive. This is particularly true for base models, which are usually not trained for instruction following and therefore continue to generate tokens until hitting their maximum token limit.
Due to computational constraints, we exclude seven generation-based benchmarks, Drop, CoQa, ReCoRD, bAbi, WebQA, TriviaQA and Fld-Default. Nevertheless, we retain two widely used generation-based benchmarks, GSM8K and NQ-Open, in our experiments.
We additionally excluded five benchmarks due to anomalies observed during fine-tuning: MedQA-4Options, LogiQA, Mutual, Mela-EN, and Swag. For these benchmarks, more than 20% of models showed no performance improvement after fine-tuning.
@@
We also excluded Paws-EN, as its corresponding model ranking under direct evaluation was negatively correlated (Kendall’s $\tau$ less than zero) with 23 out of 24 other benchmarks. We attribute this anomaly to the unusual prompting template used by lm-eval-harness.
A.2 Model Selection
See Table 2 for the complete list of models used in our experiments.
Table 2: Models considered, categorized by model family.
Family Model Name Suffix
Llama- [29] 3-8B, 3.1-8B, 3.2-1B, 3.2-3B, 3-8B-IT, 3.1-8B-IT, 3.2-1B-IT, 3.2-3B-IT
Qwen- [56] 1.5-0.5B, 1.5-1.8B, 1.5-4B, 1.5-7B, 1.5-14B, 2-0.5B, 2-1.5B, 2-7B, 2.5-0.5B, 2.5-1.5B, 2.5-3B, 2.5-7B, 2.5-14B, 1.5-0.5B-IT, 1.5-1.8B-IT, 1.5-4B-IT, 1.5-7B-IT, 1.5-14B-IT, 2-0.5B-IT, 2-1.5B-IT, 2-7B-IT, 2.5-0.5B-IT, 2.5-1.5B-IT, 2.5-3B-IT, 2.5-7B-IT, 2.5-14B-IT
Gemma- [71] 2B, 7B, 2-2B, 2-9B, 2B-IT, 7B-IT, 2-2B-IT, 2-9B-IT
GPT2- [57] 124M, 335M, 774M, 1.5B
Pythia- [8] 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B
Yi- [82] 6B, 9B, 6B-IT, 1.5-6B, 1.5-9B, 1.5-6B-IT, 1.5-9B-IT
A.3 Evaluation Setup

### v2

@@ 3.2 Downstream Ranking Agreement
As depicted in Figure 2, direct evaluation shows only modest ranking agreement between the 24 benchmarks, with an average Kendall’s $\tau$ of 0.52. This lack of agreement across benchmarks complicates model assessment and makes it challenging to aggregate results into a meaningful overall ranking (zhang2024inherent, ). In contrast, the train-before-test methodology leads to a substantial improvement in ranking agreement. Under this approach, 274 out of 276 benchmark pairs show higher Kendall’s $\tau$ scores, with the average $\tau$ rising from 0.52 to 0.76. This stronger consistency suggests that model potential ranking on one benchmark is likely to generalize to others, including practitioners’ own cases, which simplifies model comparison and selection. Notably, benchmarks that appeared to be outliers under direct evaluation, such as NQ-Open and MRPC, demonstrate much greater ranking consistency under train-before-test. For example, the average Kendall’s $\tau$ between NQ-Open and all other benchmarks improves from 0.23 to 0.74.
We further split all benchmarks into six categories (e.g., language understanding, math), see Table 1. For each category pair, we report in Figure 3 the intra-category average ranking correlations and inter-category average ranking correlations across all relevant benchmark pairs. Consistent with our previous findings, we observe reasonably poor ranking agreements across categories under direct evaluation. While one might expect high intra-category agreement—after all, tasks within the same category tend to be relatively similar—direct evaluation results in low intra-category agreement in many cases. For example, the intra-category mean Kendall’s $\tau$ is 0.54 for language understanding and 0.55 for math. This further underscores the difficulty of selecting models based on direct evaluation. Even if the goal is to choose a model that excels within a specific domain, the low intra-category agreement makes this decision challenging.
In contrast, train-before-test boosts both intra- and inter-category consistency. For example, the intra-category mean Kendall’s $\tau$ for language understanding raises from 0.52 to 0.75, as well as from 0.55 to 0.84 for the math category. Moreover, agreement between categories is often nearly as high as agreement within categories. This suggests that models with higher potential in one domain tend to also perform well across other domains after adaptation.
@@ 3.3 Perplexity Agreement
We now study the agreement between downstream benchmark rankings and perplexity rankings on general domain corpora. To do so, we collect three corpora from Wikipedia, StackExchange, and arXiv, retaining only contents from 2025. Because all models used were released before 2025, they could not have seen these texts during pretraining. Specifically, we collect 3,366 documents for Wiki, 6,001 for StackExchange and 44,384 for arXiv. These datasets are split into training, validation, and testing sets, in an 8:1:1 ratio.
@@
Figure 4: Ranking agreement between perplexity rankings and downstream benchmark rankings under direct evaluation (top) and train-before-test (bottom). Perplexity rankings are consistent with each other under both evaluation schemes, with an average Kendall’s $\tau$ of 0.76 and 0.78, respectively. However, for direct evaluation, agreement between perplexity rankings and downstream rankings is low, with an average Kendall’s $\tau$ of just 0.48. Fortunately, train-before-test results in much higher agreement between perplexity and downstream evaluations, increasing average Kendall’s $\tau$ to 0.74.
@@
Figure 5: Ranking agreement between perplexity rankings before fine-tuning (direct evaluation) and downstream benchmark rankings after fine-tuning (train-before-test) for base models (top) and instruction-tuned models (bottom). Unlike Figure 4 where both rankings in each comparison use the same evaluation scheme, here we test whether pre-fine-tuning perplexity can predict post-fine-tuning downstream performance. Base models show strong correlation (average Kendall’s $\tau$ = 0.78), suggesting perplexity is a good predictor of model potential. Instruction-tuned models show much weaker correlation (average Kendall’s $\tau$ = 0.51).
@@
We measure perplexity in bits per byte, using the lm-eval-harness library. We then compute models rankings based on the perplexity evaluations, and compare the rankings with those of the downstream benchmarks considered in earlier sections. We exclude the four Gemma models from these results, as lm-eval-harness provides unreliable perplexity measurements for Gemma models due to its rolling window implementation. See Appendix B.2 for details.
@@
The results are presented in Figure 4. In contrast to downstream tasks, perplexity rankings demonstrate strong agreement both under direct evaluation and train-before-test. Specifically, the average Kendall’s $\tau$ between the perplexity rankings is 0.76 for direct evaluation and 0.78 for train-before-test. We hypothesize that this reasonably strong agreement arises due to the smooth relationship between perplexity evaluations (brandfonbrener2024loss, ; mayilvahananllms, ).
@@
When comparing ranking agreement between perplexity evaluations and downstream benchmarks, we find that agreement is low under direct evaluation, with a mean Kendall’s $\tau$ of 0.48. This lack of agreement is concerning, as it signals a disconnect between the language modelling pre-training objective and downstream benchmark performance. Fortunately, we find that our train-before-test methodology improves ranking agreement substantially, with the mean Kendall’s $\tau$ ranking correlation between perplexity rankings and benchmark rankings rising to 0.74. This finding is reassuring: a light amount of fine-tuning on task data is sufficient to align the language modeling training objective with downstream performance. Moreover, we find that ranking agreement between perplexity and downstream evaluations is roughly similar to agreement across downstream evaluations. This suggests that, despite perplexity typically not being used for benchmarking purposes, it can be as effective a ranking metric as benchmark evaluations.
@@
Drawing inspiration from prior work (liu2023same, ; xia2023training, ; Gadre2024LanguageMS, ; benchmarkloss, ; Zhang2024WhenSM, ), we further examine the correlation between model rankings according to average perplexity across the three text corpora and average downstream performance across the 24 benchmarks. Zhang and Hardt (2024, (Gadre2024LanguageMS, )) show that, when models are trained on the same pretraining data, perplexity is well-correlated with aggregate benchmark performance. Our setup differs in that we consider a diverse set of model families, each trained on different pretraining data. Under direct evaluation, we find that the ranking correlation is modest, with a Kendall’s $\tau$ of only 0.55. We hypothesize that this relatively weak agreement is due to differences in pretraining data and instruction tuning, resulting in varying levels of exposure to benchmark tasks during training (dominguez2024training, ). Fortunately, when applying our train-before-test methodology, the ranking correlation between average perplexity and average downstream performance improves substantially, with Kendall’s $\tau$ increasing from 0.55 to 0.84.
@@
We additionally examine the agreement between perplexity prior to fine-tuning and downstream task performance after fine-tuning. That is, between direct evaluation perplexity rankings and train-before-test downstream performance rankings. We plot such ranking agreement in Figure 5, dividing models into base models and instruction-tuned models. For base models, perplexity prior to fine-tuning is a strong indicator of model potential on downstream tasks, with an average Kendall’s $\tau$ of 0.78. This indicates that, for base models, direct evaluation of perplexity is already a reasonably reliable metric for ranking models. Moreover, it indicates that the ranking consistency we observe reflects inherent model potential rather than artifacts introduced by fine-tuning.
However, the same does not hold for instruction-tuned models (average Kendall’s $\tau$ = 0.51). Instruction-tuning renders perplexity rankings unreliable, as ranking agreement is low across the board. This is to be expected: instruction fine-tuning tends to increase both benchmark performance ($\uparrow$) and perplexity ($\downarrow$) on general text corpora, thus clouding the relationship between perplexity and downstream evaluations. Fortunately, as shown earlier, train-before-test restores high ranking agreement between perplexity evaluations and downstream performance.
@@
Figure 6: Explained variance ratios of the top five principal components of the benchmark score matrix, under direct evaluation (left) and train-before-test (right). Train-before-test substantially increases the amount of variance explained by the first principal component, from 70% to 86%
@@ 3.4 Low-Ranked Model Score Matrix
So far, we have shown that comparing model potential using the train-before-test yields consistent rankings across benchmarks. We now examine the implications of this finding by analyzing the resulting matrix of model scores, where each entry $(i,j)$ corresponds to the performance of model $j$ on a benchmark $i$. We use Principal Component Analysis (PCA) to examine the structure of the model score matrix.
@@
Figure 6 shows the explained variance ratios of the first five principal components. These results support previous findings that the score matrix is of low rank (Ruan2024ObservationalSL, ). Under direct evaluation, the first five components account for 91% of the total variance. A similar trend is observed for train-before-test scores, where the first five components explain 97% of the variance. Notably, under train-before-test, the first principal component (PC1) captures a much larger share of the variance: 86%, compared to 70% for direct evaluation.
Prior works interpret PC1 scores under direct evaluation as an indication of general capability, with later principal components denoting domain-specific capabilities not captured by PC1 (e.g., reasoning ability, coding ability) (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ). Unlike out-of-the-box performance, which is controlled by multiple factors (Ruan2024ObservationalSL, ; Burnell2023RevealingTS, ), model potential is dominated by one single principal axis. It is of no surprise that PC1 also positively correlates with pre-training compute, as shown in Figure 7222We only plot models whose number of training tokens is publicly available. See Table 4 for details., which have been identified as crucial to model performances (kaplan2020scaling, ; Hoffmann2022TrainingCL, ). See detailed PC1 scores in Figure 11.
@@
We repeat our PCA analysis on the score matrix containing only Qwen models, depicted in Figure 8. Remarkably, we find that PC1 for train-before-test explains 93% of the variance, roughly as much as the variance explained by the top five principal components under direct evaluation. That is, whereas for direct evaluation the score matrix is low-rank; train-before-test renders the score matrix essentially rank one.
@@ 4 Discussion, limitations, and conclusion
Train-before-test fundamentally reframes how we interpret model evaluation. Whereas direct evaluation yields benchmark-specific rankings that often contradict one another, train-before-test harmonizes rankings across a wide array of tasks and datasets. This shift from measuring out-of-the-box performance to comparing achievable potential equips the community with a more stable and externally valid evaluation methodology.
This emphasis on model potential is particularly valuable for scenarios involving model development and adaptation. Practitioners frequently need to make decisions during model development—selecting checkpoints mid–pre-training or choosing a base model for further instruction tuning or domain-specific adaptation. In these scenarios, direct evaluation, while useful for assessing deployment readiness, is of limited relevance and utility. A model that performs poorly on direct evaluation might excel when adapted to new tasks. Train-before-test, by contrast, shows that rankings on any task will also generalize to others, offering more promising guidance for model selection.
One might argue that ranking consistency is unnecessary if we can simply choose benchmarks close to a given downstream application. However, our findings highlight three challenges with that view. First, even benchmarks that purport to measure the same skill (e.g., question answering) produce contradictory rankings under direct evaluation. Second, no benchmark perfectly captures the specifics of an application, making some degree of generalization unavoidable. Third, in real deployments, models are often adapted to varying degrees, making the potential the relevant signal for comparison.
Limitations. Train-before-test requires fine-tuning models on task-specific data before evaluation, which certainly increases the evaluation cost. However, this investment yields dividends through improved reliability. Our findings suggest that fewer benchmarks suffice under train-before-test, as rankings from one benchmark reliably transfer to others. This reduction in required evaluations can offset the per-benchmark cost increase. A second problem is that, unfortunately, many benchmarks no longer come with training data, making it more difficult to apply train-before-test. In light of our findings, we recommend that future benchmarks provide fine-tuning data for the benchmark. A third limitation is that some commercial model providers do not easily allow fine-tuning of their models. We contend that in this case the problem is with the model provider. There is clearly scientific value in creating an ecosystem of models that can be fine-tuned. Train-before-test evaluation creates additional incentives for making models easy to fine-tune.
Conclusion. Overall, train-before-test complements existing evaluation practices by distinguishing between performance and potential. cImportantly, potential comparison is not intended to replace direct evaluation—both serve distinct purposes. Direct evaluation remains useful for understanding immediate deployment readiness, while potential comparison provides insights into adaptability and development prospects. Together, they offer a more complete picture of model capabilities. We believe that adopting train-before-test as a standard alongside direct evaluation can significantly improve the reliability, interpretability, and practical utility of the model evaluation ecosystem.
@@ A.1 Benchmark Selection
We begin our study with the lm-eval-harness package [26], which offers a comprehensive suite of language model benchmarks. To accommodate the train-before-test methodology which requires a dedicated training set for fine-tuning, we first identify benchmarks that provide at least 1,000 training examples. This yields a total of 37 benchmarks.
@@
These benchmarks can be broadly categorized as 28 likelihood-based and 9 generation-based benchmarks. Likelihood-based evaluations test for the likelihood of different completions given some input string; for example, different answer choices given a multiple-choice input question. Since the number of completions is usually small, likelihood-based evaluations are generally compute-efficient.
Generation-based evaluations, in contrast, generate some output response given an input query. If responses tend to be long, then generation-based evaluations naturally become compute-intensive. This is particularly true for base models, which are usually not trained for instruction following and therefore continue to generate tokens until hitting their maximum token limit. These generation-based benchmarks are also over-challenging for smaller models with limited parameters, such as GPT-2 [64]. Therefore, we exclude seven generation-based benchmarks, Drop, CoQa, ReCoRD, bAbi, WebQA, TriviaQA and Fld-Default. Nevertheless, we retain two widely used generation-based benchmarks, GSM8K and NQ-Open, in our experiments.
@@
We additionally excluded five benchmarks due to anomalies observed during fine-tuning: MedQA-4Options, LogiQA, Mutual, Mela-EN, and Swag. For these benchmarks, more than 20% of models, most of which are small models with fewer than 0.5B parameters, showed no performance improvement after fine-tuning. We also excluded Paws-EN, as its corresponding model ranking under direct evaluation was negatively correlated (Kendall’s $\tau$ less than zero) with 23 out of 24 other benchmarks. We attribute this anomaly to the unusual prompting template used by lm-eval-harness.
Our final selection includes 24 benchmarks: MNLI [85], QNLI [65], RTE [17, 30, 7], CoLA [81], SST-2 [73], MRPC [19], QQP, WiC [63], ANLI [59], Winogrande [46], CommonsenseQA [76], Hellaswag [90], Social-IQA [71], OpenBookQA [57], NQ-Open [44], BoolQ [14], ARC-Easy, ARC-Challenge [15], SciQ [84], PIQA [9], MathQA [3], GSM8K [16], MedMCQA [62], HeadQA [80].
A.2 Evaluation Setup